OpenStudy (anonymous):

(4 1/3)/(-1 1/3) ^ ca n sumone help me solve fastt

OpenStudy (anonymous):

3 *4 =12 +1, 13/3 3*-1=-3+1=-2/3 (13/3)/(-2/3)=-13/2

OpenStudy (anonymous):

when dividing, multiply the reciprocal

OpenStudy (lalaly):

\[\frac{13}{3} \times \frac{3}{-2} = \frac{13}{-2}\]